OpenStudy (anonymous):

Use the formula A=P(l+rt) to find the value od P when A=2912, r=0.06, and t=2 P=

OpenStudy (mathmate):

A=P(l+rt) so evaluate: P=A/(1+rt)

OpenStudy (anonymous):

I need to put it as an integer or a decimael rounded to the nearest hundredth as needed

OpenStudy (mathmate):

You'd use the calculator to evaluate P, then you do the rounding afterwards.

OpenStudy (anonymous):

Sorry I dont have a calculator

OpenStudy (mathmate):

You have two choices: 1. type in the expression into google, which should give you the answer 2. use one of the many online calculators 3. practise your long division and do it on paper.

OpenStudy (mathmate):

sorry, three choices

OpenStudy (anonymous):

ok thank u I am 55 years old trying to get through this math course and it is really stressing me out but thank u for all ur help

OpenStudy (mathmate):

If you do not have a calculator, would you be able to take your exam?

OpenStudy (anonymous):

not really

OpenStudy (mathmate):

Have you tried using Google as a calculator? Just have to type in 54*3 it will return 162 as the answer, and not a search.